# supplement_jasa

### v2

@@ Abstract
Building artificially intelligent geospatial systems require rapid delivery of spatial data analysis at massive scales with minimal human intervention. Depending upon their intended use, data analysis may also entail model assessment and uncertainty quantification. This article devises transfer learning frameworks for deployment in artificially intelligent systems, where a massive data set is split into smaller data sets that stream into the analytical framework to propagate learning and assimilate inference for the entire data set. Specifically, we introduce Bayesian predictive stacking for multivariate spatial data and demonstrate rapid and automated analysis of massive data sets. Furthermore, inference is delivered without human intervention without excessively demanding hardware settings. We illustrate the effectiveness of our approach through extensive simulation experiments and in producing inference from massive datasets on sea surface temperatures and vegetation index that are indistinguishable from traditional (and more expensive) statistical approaches.
@@ 1 Introduction
Geospatial artificial intelligence (GeoAI) is a rapidly evolving discipline at the interface of machine learning and spatial data science that attempts to harness the analytical capabilities of Artificial Intelligence (AI) to analyze massive amounts of geographic data and conduct data driven scientific discoveries in diverse domains within the environmental and physical sciences. The area, while still fledgling, presents opportunities for statisticians to devise essential data analytic tools that should comprise an artificially intelligent “geospatial” system. Constructing comprehensive AI-based geospatial systems involve expertise in computer science relating to efficiently designed Geographic Information Systems (GIS) databases, conventional machine learning tools based upon algorithms for spatial interpolation, and statistical or probabilistic learning tools to offer scientific inference. This manuscript concerns itself primarily with statistical learning tools for artificially intelligent GeoAI systems.
@@
What, if at all, should be the role of formal statistical inference in GeoAI? Formal inference for spatial and spatial-temporal random fields enjoy a venerable intellectual presence spanning several decades of theoretical developments within classical and Bayesian paradigms (see, e.g., Cressie,, 1993; Stein,, 1999; Gelfand et al.,, 2010; Cressie and Wikle,, 2011; Banerjee et al.,, 2015). With increasing complexity in spatially oriented datasets, statisticians have built richly structured hierarchical models in their pursuit of “full” probabilistic inference. The term “full” loosely refers to approaches that attempt to estimate all unknowns in the model including all parameters (irrespective of how well they are identified by the data), random effects or partial realizations of the process, and predictive random variables for probabilistic interpolation at arbitrary locations (and time points in spatial-temporal data).
Statisticians would, we believe, broadly agree that GeoAI offers sufficient intellectual space to accommodate rigorous probabilistic learning as a crucial constituent in artificially intelligent data analysis systems. Spatial modeling relies upon Gaussian processes (GPs) to model dependence and achieves superior predictive inference. While GPs afford flexibility and are a conspicuous choice in spatial modeling, their covariance kernels do not necessarily yield computationally exploitable structures for the covariance matrices and the computational requirements for full inference become onerous and impracticable for massive datasets.
@@
This presents a conundrum for statisticians: should we strive to completely retain the probabilistic rigors of theoretical inference, which may be computationally burdensome, or even impracticable, given the massive amounts of data, or should we concede uncertainty quantification and adopt algorithm-driven data analysis that would easily scale massive data sets? Spatial data analysis at unprecedented scales requires some concessions from conventional decision-theoretic paradigms as has been reflected, for example, in a recent comprehensive case study by Zhou et al., (2022), who synthesize statistical models with machine learning algorithms to measure temporal trends and spatial distribution of housing vitality by exploiting information from multiple sources (also see the discussion by Banerjee,, 2022, for conventional Bayesian modeling perspectives and its challenges). Other efforts in melding machine learning methods with formal spatial statistics include deep learning for spatial data (Zammit-Mangion et al.,, 2022; Wikle and Zammit-Mangion,, 2023, with the latter offering a comprehensive review of spatial machine learning) and, in a very related domain, for deep GP emulation in computer experiments (Sauer et al., 2023b, ), and spatial random forests (Georganos et al.,, 2021; Talebi et al.,, 2022; Saha et al.,, 2023).
@@
Even a cursory review reveals a significant literature on statistical methods for massive spatial datasets, which is too vast to be summarized here (see, e.g., Banerjee,, 2017; Heaton et al.,, 2017, and references therein). Within the hierarchical setup $[data\mid process]\times[process\mid parameters]\times[parameters]$, inference proceeds from spatial processes that scale massive data sets. Examples range from reduced-rank processes or subsets of regression approaches (Quiñonero-Candela and Rasmussen,, 2005; Cressie and Johannesson,, 2008; Banerjee et al.,, 2008; Wikle,, 2010, and references therein), multi-resolution approaches (Nychka et al.,, 2015; Katzfuss,, 2017), and graph-based models Vecchia, (1988); Datta et al., (2016); Katzfuss and Guinness, (2021); Dey et al., (2022); Sauer et al., 2023a . Full inference typically requires efficient Markov chain Monte Carlo (mcmc) (Finley et al.,, 2019), variational approximations (Ren et al.,, 2011; Wu et al.,, 2022; Cao et al.,, 2023) and a significant body of literature on Gaussian Markov random field approximations (Rue et al.,, 2009; Lindgren et al.,, 2011, and references therein) in conjunction with integrated nested Laplace approximations (inla) for computing the marginal distributions of the process at given locations.
@@
The aforementioned methods focus upon richness of statistical inference, but almost invariably involve a significant amount of human intervention to analyze spatial data. Even the simplest of geostatistical datasets require exploratory data analysis to learn about aspects of the underlying process that are weakly identified by the data (Zhang,, 2004; Tang et al.,, 2021). Building a GeoAI system, on the other hand, will require minimizing human intervention in offering a robust framework for spatial data analysis. This presents enough challenges that preclude a comprehensive solution in its entirety within the scope of a single manuscript. Nevertheless, we devise a spatial data analytic framework that holds significant promise for GeoAI. The premise of this approach relies upon two basic tenets: (i) model-based statistical inference for underlying spatial processes (including multivariate processes) in a robust and largely automated manner with minimal human input; and (ii) achieving such inference for truly massive amounts of data without resorting to iterative algorithms that may require significant human intervention to diagnose convergence (such as in mcmc).
@@
We wish to retain the benefits of Bayesian hierarchical models, convincingly documented in the aforementioned manuscripts, with regard to spatial interpolation of the underlying stochastic process. In particular, we seek predictive inference over the uncountable collection of points in the spatial domain while propagating uncertainty about the model’s estimation in such inference. However, some compromises with regard to richness of models are inevitable from the standpoint of GeoAI. At the same time, statistical distribution theory has much to offer and we harness analytical closed-form distribution theory within a family of matrix-variate distributions to deliver inference. This is possible only by “fixing” certain spatial correlation kernel parameters that are weakly identified and hinder automated inference. Our approach relies upon carrying out fast and exact inference based upon the matrix-variate distributions and then assimilating the inference using Bayesian predictive stacking or bps (Wolpert,, 1992; Yao et al.,, 2018). Stacking (Breiman,, 1996) is primarily related to predictions by minimizing the generalization error rate of one or more models.
@@
To analyze massive amounts of data to be accessed in a GeoAI system, we devise Bayesian transfer learning over a streaming set of spatial datasets. We envisage a setting where the data stream into the GeoAI system in the form of $K$ subsets, where each subset is composed of $\approx n/K$ locations with $K$ being a function of computational resources available to GeoAI. We analyze and transfer learning from one dataset to the next and assimilate inference from these subsets to the full data set over $n$ locations. The concept of delegating Bayesian calculations to a group of independent datasets is intuitive and has been explored from diverse perspectives. Examples include Consensus Monte Carlo (cmc) (Scott et al.,, 2016) and its kernel-based adaptations (Rendell et al.,, 2021), Bayesian meta-analysis in clinical applications (Parmigiani,, 2002, see, e.g., Chapter 4 in) and in spatially-temporally structured data (Bell et al.,, 2005; Kang et al.,, 2011); and, more recently, approaches such as “meta-kriging” using geometric medians of posterior distributions (Srivastava et al.,, 2015; Minsker et al.,, 2017; Guhaniyogi and Banerjee,, 2018, 2019; Guhaniyogi et al.,, 2019). However, each of these approaches have, hitherto, relied upon MCMC or other iterative algorithms for estimating the weakly identifiable parameter. We, instead, propose to execute transfer learning using a double stacking approach. We first obtain stacked inference for each of the parameters within each subset and, then, assimilate inference across subsets by a second stacking algorithm. A flowchart is supplied in Figure 1 comprising three steps: (i) partition data into smaller subsets; (ii) analyze each subset by stacking conjugate posteriors in closed form; and (iii) stack across data sets to deliver inference for the entire data.
@@
The contents of the paper evolve in accordance with Figure 1. Section 2 outlines Bayesian transfer learning using double bps. Section 3 provides details on computational aspects. Section 4 presents simulation experiments designed for empirically evaluating double bps. Section 5 analyzes a very large climate data set with observed locations in the millions at global scales. Finally, Section 6 concludes with some discussion. References to an accompanying Supplement are provided throughout the article and after Section 6.
@@ 2 Bayesian transfer learning
Transfer learning (TL) broadly refers to propagating knowledge from one task to accomplish a different task. We consider transferring inference from one subset to the next in a stream of subsets to analyze the entire dataset. Learning from spatial random fields is challenging due to inherent dependencies in the data and we seek to infer about a latent spatial process (uncountable collection of unobserved random variables) with limited data (partially observed finite realization of the process). We develop an automated approach for GeoAI.
@@ 2.1 Divide-and-Conquer Multivariate Bayesian inference
Spatial random fields immediately present a challenge. The above method delivers full inference without any loss of information only if $Y_{k}$’s are reasonably assumed to be independent (or exchangeable) across blocks. If each row of $Y$ corresponds to one of $n$ spatial locations so that $V$ is an $n\times n$ spatial correlation matrix, the above scheme requires that each $V_{k}$ is the spatial correlation matrix constructed from the $n_{k}$ spatial locations in $\mathscr{D}_{k}$. Independence across the $K$ blocks may yield reasonable inference, especially if it is possible to design the blocks such that the spatial correlation across blocks do not impact overall inference, but this will require significant human intervention or a separate AI system for designing such blocks. We seek to avoid this in GeoAI.
Instead, we devise a method for assimilating the learning from each of these blocks using predictive stacking. We exploit the fact that $V$ is indexed by a small number of parameters in a spatial correlation kernel. Fixing these parameters fixes $V$, and hence $V_{k}$ for each $k$, yielding closed-form posterior inference on $\beta$ and $\Sigma$ based upon the entire dataset as described above. Stacking combines these analytically accessible distributions using an optimal set of weights that are computed using a convex optimization algorithm. These weights are then used to reconstruct the posterior and predictive distributions for the spatial random field without imposing block independence. It is worth remarking here that existing “meta-kriging” and related “divide and conquer” approaches (e.g., Guhaniyogi and Banerjee,, 2018; Scott et al.,, 2016, and other references on “divide and conquer” methods provided in Section 1) analyze subsets of data using mcmc, which is expensive and not fully automated. A key distinction of the current manuscript is that we abandon all iterative estimation algorithms, let alone expensive mcmc, and focus on assimilating inference using closed forms of distributions. The matrix-variate formulation we pursue enhances flexibility and scope of applications.
@@ 2.2 Bayesian stacking of predictive densities
where $\mathscr{D}_{-i}$ is the dataset excluding the $i$-th observation, or block (indexed by a row) of observations in the matrix $Y$, and $\mathscr{M}=\left(\mathscr{M}_{1},\ldots,\mathscr{M}_{J}\right)$ are $J$ different models. For any given dataset $\mathscr{D}$, these $J$ different models correspond to fixed spatial correlation kernel parameters in $V$ and deliver closed-form analytic distributions. Solving (2.3) minimizes the Kullback-Leibler divergence from the true predictive distribution and is executed using convex optimization (Grant and Boyd,, 2008; Fu et al.,, 2020). While the true predictive distribution is unknown, we can use a leave-one-out (loo) estimate of the expected value of the score (see. e.g., Yao et al.,, 2018, for details), which requires fitting the model $n$ times as we exclude one row of the data $\mathscr{D}$ at a time so we prefer K-fold cross-validation as a more cost-effective method for generating predictions (Breiman,, 1996).
@@ 2.3 Accelerated learning for spatial random fields
We turn to adapting the aforementioned concepts into an automated learning framework for GeoAI. Let $\mathcal{S}=\left\{s_{1},\ldots,s_{n}\right\}\subset\mathcal{D}$ be a set of $n$ locations yielding observations on $q$ possibly correlated outcomes collected into a $q\times 1$ vector $y(s)=(y_{1}(s),\ldots,y_{q}(s))^{{\mathrm{\scriptscriptstyle T}}}$ for each $s\in\mathcal{S}$. We collect these measurements into the $n\times q$ matrix $Y=[y_{j}(s_{i})^{{\mathrm{\scriptscriptstyle T}}}]$ for $i=1,\ldots,n$ and $j=1,\ldots,q$. Let $X=[x(s_{i})^{{\mathrm{\scriptscriptstyle T}}}]$ be $n\times p$ with rows $x(s_{i})^{{\mathrm{\scriptscriptstyle T}}}$ consisting of $p<n$ explanatory variables at location $s_{i}\in\mathcal{S}$; we assume $X$ has rank $p$. We cast this into (2.2), but explicitly introduce a latent $q\times 1$ spatial process $\omega(s)$ as
@@
The covariance among the elements of $y(s)$ within $s$ is given by the elements of $\left(\alpha^{-1}-1\right)\Sigma$. Thus, $\Sigma$ is the within-location (nonspatial) dependence among the outcomes adjusted by a scale of $\left(\alpha^{-1}-1\right)$ to accommodate additional variation at local scales. The parameter $\alpha$ is the ratio of the “partial sill” to the “sill” in classical geostatistics. In the special case when $\Sigma=\sigma^{2}\mathbb{I}_{q}$, we interpret $\alpha=\sigma^{2}/\left(\sigma^{2}+\tau^{2}\right)$, where $\tau^{2}$ is the measurement error variance (“nugget”) and $\alpha$ is the ratio of the spatial variance (partial sill) to the total variance (sill).
@@
where $V^{\star}_{\gamma}=\left[\begin{array}[]{cc}\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}X+M_{0}^{-1}&\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}\\ \frac{\alpha}{1-\alpha}X&\rho_{\phi}^{-1}(\mathcal{S},\mathcal{S})+\frac{\alpha}{1-\alpha}\mathbb{I}_{n}\end{array}\right]^{-1}$ and $\mu_{\gamma}^{\star}=V^{\star}_{\gamma}\left[\begin{array}[]{c}\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}Y+m_{0}\\ \frac{\alpha}{1-\alpha}Y\end{array}\right]$, $\Psi^{\star}=\Psi_{0}+\frac{\alpha}{1-\alpha}Y^{{\mathrm{\scriptscriptstyle T}}}Y+m_{0}^{{\mathrm{\scriptscriptstyle T}}}M_{0}m_{0}-\mu_{\gamma}^{*{\mathrm{\scriptscriptstyle T}}}V^{\star-1}_{\gamma}\mu_{\gamma}^{\star}$ and $\nu^{\star}=\nu_{0}+n$.

### v3

# Bayesian Transfer Learning for Artificially Intelligent Geospatial Systems: A Predictive Stacking Approach

@@ Abstract
Building artificially intelligent geospatial systems requires rapid delivery of spatial data analysis on massive scales with minimal human intervention. Depending upon their intended use, data analysis can also involve model assessment and uncertainty quantification. This article devises transfer learning frameworks for deployment in artificially intelligent systems, where a massive data set is split into smaller data sets that stream into the analytical framework to propagate learning and assimilate inference for the entire data set. Specifically, we introduce Bayesian predictive stacking for multivariate spatial data and demonstrate rapid and automated analysis of massive data sets. Furthermore, inference is delivered without human intervention without excessively demanding hardware settings. We illustrate the effectiveness of our approach through extensive simulation experiments and in producing inference from massive dataset on vegetation index that are indistinguishable from traditional (and more expensive) statistical approaches.
@@ 1 Introduction
Geospatial artificial intelligence (GeoAI) is a rapidly evolving discipline at the interface of machine learning and spatial data science that attempts to harness the analytical capabilities of Artificial Intelligence (AI) to analyze massive amounts of geographic data for data driven scientific discoveries in the environmental and physical sciences. The area, while still fledgling, presents opportunities for statisticians to devise essential data analytic tools that should comprise an artificially intelligent “geospatial” system. This manuscript concerns itself primarily with statistical learning tools for artificially intelligent GeoAI systems.
@@
What should be the role of statistical inference in GeoAI? Formal inference for spatial random fields enjoy a venerable intellectual presence spanning several decades of theoretical developments within classical and Bayesian paradigms (see, e.g., Cressie, , 1993; Stein, , 1999; Gelfand et al., , 2010; Cressie and Wikle, , 2011; Banerjee et al., , 2015). Statisticians have built richly structured hierarchical models in pursuit of “full” probabilistic inference for richly structured spatial data. The term “full” loosely refers to approaches that estimate all unknowns in the model including all parameters (irrespective of how well they are informed by the data), random effects, and predictive random variables for probabilistic interpolation at arbitrary locations (and time points in spatial-temporal data).
@@
Spatial modeling relies upon Gaussian processes (GPs) to model dependence and achieves superior predictive inference. While GPs afford flexibility and are a conspicuous choice in spatial modeling, they do not generally offer computationally exploitable structures for covariance matrices. The computational requirements for full inference become onerous and impracticable for massive datasets. This presents a conundrum: should we strive to completely retain the probabilistic rigors of theoretical inference, which may be computationally burdensome, or even impracticable, given the massive amounts of data, or should we concede uncertainty quantification and adopt algorithm-driven data analysis that would easily scale massive data sets? Spatial data analysis at unprecedented scales requires some concessions from conventional decision-theoretic paradigms as has been reflected, for example, in a recent comprehensive case study by Zhou et al., (2022), who synthesize statistical models with machine learning algorithms to measure temporal trends and spatial distribution of housing vitality by exploiting information from multiple sources (also see the discussion by Banerjee, , 2022, for conventional Bayesian modeling perspectives and its challenges). Other efforts in melding machine learning methods with formal spatial statistics include deep learning for spatial data (Zammit-Mangion et al., , 2022; Wikle and Zammit-Mangion, , 2023, with the latter offering a comprehensive review of spatial machine learning) and, in a very related domain, for deep GP emulation in computer experiments (Sauer et al., 2023b, ), and spatial random forests (Georganos et al., , 2021; Talebi et al., , 2022; Saha et al., , 2023).
@@
Even a cursory review reveals a significant literature on methods for massive spatial datasets, which is too vast to be summarized here (see, e.g., Banerjee, , 2017; Heaton et al., , 2017). In hierarchical models $data\mid process]\times[process\mid parameters]\times[parameters]$, inference proceeds from spatial processes that scale massive data sets. Examples range from reduced-rank or subsets of regression approaches (see, e.g., Quiñonero-Candela and Rasmussen, , 2005; Cressie and Johannesson, , 2008; Banerjee et al., , 2008; Wikle, , 2010), multi-resolution approaches (Nychka et al., , 2015; Katzfuss, , 2017), and graph-based models Vecchia, (1988); Datta et al., (2016); Katzfuss and Guinness, (2021); Dey et al., (2022); Sauer et al., 2023a . Full inference typically requires Markov chain Monte Carlo (mcmc) (Finley et al., , 2019), variational approximations (Ren et al., , 2011; Wu et al., , 2022; Cao et al., , 2023) and a significant body of literature on Gaussian Markov random field approximations (Rue et al., , 2009; Lindgren et al., , 2011) in conjunction with integrated nested Laplace approximations (inla) for computing the marginal distributions of the process at given locations.
@@
The aforementioned methods focus upon richness of statistical inference, but almost invariably involve a significant amount of human intervention to analyze spatial data. Even the simplest geostatistical data require exploratory data analysis to learn about aspects of the underlying process that are weakly identified by the data (Zhang, , 2004; Tang et al., , 2021). Building a GeoAI system, on the other hand, will require minimizing human intervention in offering a robust framework for spatial data analysis. This presents enough challenges that preclude a comprehensive solution in its entirety within the scope of a single manuscript. Nevertheless, we devise a spatial data analytic framework that holds significant promise for GeoAI. The premise of this approach relies upon two basic tenets: (i) model-based statistical inference for underlying spatial processes (including multivariate processes) in a robust and largely automated manner with minimal human input; and (ii) achieving such inference for truly massive amounts of data without resorting to iterative algorithms that may require significant human intervention to diagnose convergence (such as in mcmc).
@@
We retain the benefits of posterior learning of the underlying stochastic process. In particular, we obtain predictive inference over the uncountable collection of points in the spatial domain while propagating uncertainty about the model’s estimation in such inference. However, some compromises with regard to richness of models are inevitable from the standpoint of GeoAI. At the same time, statistical distribution theory has much to offer and we harness analytical closed-form distribution theory within a family of matrix-variate distributions to deliver inference. This is possible only by “fixing” certain spatial correlation kernel parameters that are weakly identified and hinder automated inference. We execute fast and exact inference based upon the matrix-variate distributions and then assimilate the inference using Bayesian predictive stacking or bps (Wolpert, , 1992; Yao et al., , 2018). Stacking (Breiman, , 1996) is primarily related to predictions by minimizing the generalization error rate of one or more models. Bayesian predictive stacking (bps) is a particular type of Bayesian predictive synthesis (see, e.g., Yao et al., , 2018; McAlinn and West, , 2019; McAlinn et al., , 2020; Tallman and West, , 2023; Cabel et al., , 2025, for different perspectives).
@@
A GeoAI system will access massive amounts of data. We devise Bayesian transfer learning over a streaming set of spatial datasets. The data streams into the GeoAI system in the form of $K$ subsets, each with $\approx n/K$ locations. We analyze and propagate inference from one subset to the next and assimilate inference from these subsets. Delegating Bayesian calculations to a group of independent datasets is intuitive and has been explored from diverse perspectives. Examples include Consensus Monte Carlo (cmc) (Scott et al., , 2016) and its kernel-based adaptations (Rendell et al., , 2021), Bayesian meta-analysis in clinical applications (see, e.g., Chapter 4 in Parmigiani, , 2002) and in spatially-temporally structured data (Bell et al., , 2005; Kang et al., , 2011); and, more recently, “meta” approaches using diverse distributed computing architectures (Srivastava et al., , 2015; Deisenroth and Ng, , 2015; Minsker et al., , 2017; Srivastava et al., , 2018; Guhaniyogi and Banerjee, , 2018, 2019; Guhaniyogi et al., , 2019, 2022, 2023). Most approaches have relied on iterative algorithms for estimating weakly identifiable parameters. We propose transfer learning using a double stacking approach. We first obtain stacked inference for each of the parameters within each subset and, then, assimilate inference across subsets by a second stacking algorithm. A flowchart is supplied in Figure 1 comprising three steps: (i) partition data into smaller subsets; (ii) analyze each subset by stacking conjugate posteriors in closed form; and (iii) stack across data sets to deliver inference for the entire data.
@@
This article evolves according to Figure 1. Section 2 outlines Bayesian transfer learning using double bps (dbps), Section 3 provides details on computation, Section 4 presents simulation experiments to illustrate and evaluate double bps and its applicability to amortized inference. Section 5 analyzes a vegetation index data set with observed locations in the millions on global scales. Finally, Section 6 concludes with some discussion. References to an accompanying Supplement are provided throughout the article and after Section 6.
@@ 2 Bayesian transfer learning
Transfer learning (TL) broadly refers to propagating knowledge from one task to accomplish a different task (see Suder et al., , 2023, and references therein). We devise transferring inference from one subset to the next in a stream of subsets to analyze the entire dataset. Learning from spatial random fields is challenging because of limited information available for the latent spatial process (uncountable collection of unobserved random variables) using a partially observed finite realization. We develop an automated approach for GeoAI.
@@ 2.1 Divide-and-Conquer Multivariate Bayesian inference
Spatial random fields immediately present a challenge. The above method delivers inference without loss of information only if $Y_{k}$’s are independent between blocks. If each row of $Y$ corresponds to a spatial location so $V$ is an $n\times n$ spatial correlation matrix, then each $V_{k}$ is the spatial correlation matrix constructed from $n_{k}$ spatial locations in $\mathscr{D}_{k}$. The independence among $K$ blocks may yield reasonable inference if we can design the blocks such that the spatial correlation between blocks does not impact the general inference. However, designing such blocks requires significant human intervention. We seek to avoid this in GeoAI.
Instead, we assimilate statistical learning from each of these blocks using predictive stacking. We exploit the fact that $V$ is indexed by a small number of parameters in a spatial correlation kernel. Fixing these parameters fixes $V$, and hence $V_{k}$ for each $k$, producing closed-form posterior inference on $\beta$ and $\Sigma$ based on the entire dataset as described above. Stacking combines these analytically accessible distributions using an optimal set of weights that are computed using a convex optimization algorithm. These weights are then used to reconstruct the posterior and predictive distributions for the spatial random field without imposing block independence. Existing “meta-kriging” and related “divide and conquer” approaches (e.g., Guhaniyogi and Banerjee, , 2018; Scott et al., , 2016, and other references on “divide and conquer” methods provided in Section 1) analyze subsets of data using mcmc, which is expensive and not fully automated. A key distinction of the current manuscript is that we abandon all iterative estimation algorithms, let alone expensive mcmc, and focus on assimilating inference using closed forms of distributions.
@@ 2.2 Bayesian stacking of predictive densities
where $\mathscr{D}_{-i}$ is the dataset excluding the $i$-th block (indexed by a row) of observations in $Y$, and $\mathscr{M}=\left(\mathscr{M}_{1},\ldots,\mathscr{M}_{J}\right)$ are $J$ different models. For any given dataset $\mathscr{D}$, these $J$ different models correspond to fixed spatial correlation kernel parameters in $V$ and deliver closed-form analytic distributions. Solving (2.3) minimizes the Kullback-Leibler divergence from the true predictive distribution using convex optimization (Grant and Boyd, , 2008; Fu et al., , 2020). While the true predictive distribution is unknown, we use a leave-one-out (loo) estimate of the expected value of the score (see. e.g., Yao et al., , 2018, for details), which requires fitting the model $n$ times as we exclude one row of the data $\mathscr{D}$ at a time. We prefer K-fold cross-validation as a more cost-effective method for generating predictions (Breiman, , 1996).
@@ 2.3 Accelerated learning for spatial random fields
Let $\mathcal{S}=\left\{s_{1},\ldots,s_{n}\right\}\subset\mathcal{D}$ be a set of $n$ locations yielding observations on $q$ possibly correlated outcomes collected into a $q\times 1$ vector $y(s)=(y_{1}(s),\ldots,y_{q}(s))^{{\mathrm{\scriptscriptstyle T}}}$ for each $s\in\mathcal{S}$. We collect these measurements into the $n\times q$ matrix $Y=[y_{j}(s_{i})^{{\mathrm{\scriptscriptstyle T}}}]$ for $i=1,\ldots,n$ and $j=1,\ldots,q$. Let $X=[x(s_{i})^{{\mathrm{\scriptscriptstyle T}}}]$ be $n\times p$ with rows $x(s_{i})^{{\mathrm{\scriptscriptstyle T}}}$ consisting of $p<n$ explanatory variables at location $s_{i}\in\mathcal{S}$; we assume $X$ has rank $p$. We introduce latent spatial processes, $\omega_{j}(s)$, for each outcome $y_{j}(s)$ to capture spatial dependence and a $q\times q$ covariance matrix, $\Sigma$, to capture non-spatial dependence among the elements of $y(s)$ within $s$. This matrix is typically adjusted by a scale factor $\left(\alpha^{-1}-1\right)$ to accommodate additional variation at local scales. For example, setting $\alpha=\sigma^{2}/\left(\sigma^{2}+\tau^{2}\right)$ where $\sigma^{2}$ and $\tau^{2}$ denote variances for the spatial process and measurement error (“nugget”) implies $\alpha$ is the ratio of the spatial variance (partial sill) to the total variance (sill) gleaned from a variogram in classical geostatistics.
We cast this into (2.2), but explicitly introduce a latent $q\times 1$ spatial process $\omega(s)$ as
@@
where $V^{\star}_{\gamma}=\begin{bmatrix}\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}X+M_{0}^{-1}&\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}\\ \frac{\alpha}{1-\alpha}X&\rho_{\phi}^{-1}(\mathcal{S},\mathcal{S})+\frac{\alpha}{1-\alpha}\mathbb{I}_{n}\end{bmatrix}^{-1}$ and $\mu_{\gamma}^{\star}=V^{\star}_{\gamma}\begin{bmatrix}\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}Y+m_{0}\\ \frac{\alpha}{1-\alpha}Y\end{bmatrix}$, $\Psi^{\star}=\Psi_{0}+\frac{\alpha}{1-\alpha}Y^{{\mathrm{\scriptscriptstyle T}}}Y+m_{0}^{{\mathrm{\scriptscriptstyle T}}}M_{0}m_{0}-\mu_{\gamma}^{*{\mathrm{\scriptscriptstyle T}}}V^{\star-1}_{\gamma}\mu_{\gamma}^{\star}$ and $\nu^{\star}=\nu_{0}+n$.